OpenStudy (anonymous):

2/(a+3) - (a^2+3a-10)/(a^3+5a^2+6a)=(a-4)/(a^2+3a) Solve for a. Please explaiiin to me how this is -6?

OpenStudy (anonymous):

PLEASE, someone answer this question, outcastlilgurl has been waiting here for like an hour

OpenStudy (anonymous):

outcast, just wait a few minutes, be patient, sstarica always helps, she is like an angel

OpenStudy (anonymous):

alright, first you need to set them to the same denominator, you'll get :\[= [2a(a+2) - (a+5)(a-2) - (a-4)] / a(a+3)(a+2) \]\[= (2a^2 + 4a - a^2 - 3a + 10 - a^2 + 2a - 8)/a(a+3)(a+2)\]\[= 3(a+3)/a(a+2)(a+3) = 3/a(a+2)\] that's what I got ._.

OpenStudy (anonymous):

lol, andy

OpenStudy (anonymous):

are you sure it must be = -6?

OpenStudy (anonymous):

he got tired of waiting so he went downtown to get a coke

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

or , you move the 2nd portion to the other side and simplify to get the same denom. then cross multiply and you'll get in the end: 5a^2 + 12a - 9 = 0 then a = 3/5 and a = -3

OpenStudy (radar):

I get -6

OpenStudy (radar):

Which proved to be wrong........too many numbers Ilost ttrack.

OpenStudy (anonymous):

geez well then the math answer key must be wrong...

OpenStudy (anonymous):

and I'm a girl XD